OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (4, −7) and a directrix of y = −15. Put the equation in standard form.

OpenStudy (anonymous):

answers f(x) = one sixteenthx2 − 8x + 11 f(x) = one sixteenthx2 − 8x − 10 f(x) = one sixteenthx2 − one halfx + 11 f(x) = one sixteenthx2 − one halfx − 10

OpenStudy (amistre64):

well, we should be able to see that the vertex will be situated on the x=4 axis, with a y value that is halfway between the focus and directrix values .

OpenStudy (amistre64):

they are nice enough to give you a slope of 1/16th on every option to take care of any other computations

OpenStudy (anonymous):

and

OpenStudy (amistre64):

well if you know a vertex (a,b) and a leading coefficeint (m) the equation becomes: y = m(x-a)^2 + b which we can then expand to the option they want

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can u help me on answering a few questions please

OpenStudy (anonymous):

For this adventure, you and world renowned Professor Sherlock McMerlock are traveling to the Lost Island of Laplaya. When the boat arrives on the island shore, you and Professor McMerlock disembark on your adventure. You trudge through the jungles and arrive at three impressively large doors. About eye level on each door is an intricately carved keyhole. Directions are scratched into the wood above each keyhole. To open these doors, you must speak three functions in standard form. One function, f(x), with two real rational solutions. One function, g(x), with two real irrational solutions. One function, h(x), with two complex solutions.

OpenStudy (amistre64):

please do not post your own question in someone elses question.